OpenStudy (anonymous):

write an equation of a line in slope intercept form that passes through the points (2,4) and (5,4).

OpenStudy (anonymous):

y2-y1 over x2-x1 = slop

OpenStudy (anonymous):

Use the point-slope form y-y1 = m(x-x1) where m is the slope and (x1-y1) is any of the two points given. Firstly, find the slope.

OpenStudy (anonymous):

y2-y1/x2-x1

OpenStudy (anonymous):

4-4/5-2

OpenStudy (anonymous):

0/3 which is 0

OpenStudy (anonymous):

yes that's right @tylerj37

OpenStudy (anonymous):

slope of 0?

OpenStudy (anonymous):



OpenStudy (anonymous):

Having difficulty anywhere @tylerj37 ??